OpenStudy (anonymous):

In the equation y equals one-fourth x plus 8, what is y when x is twenty?

OpenStudy (ciarán95):

The description given in the question can be represented by the following equation: \[y = \frac{ 1 }{ 4 }x + 8\] So, if we want to find out what y is when x = 20, we substitute in 20 for x in the equation above and solve for y. \[y = \frac{ 1 }{ 4 }(20) + 8 = 5 + 8 = ?\]